Find a b and c of this ellipse 9x²-3y²-36x-6y+12=0
Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52803)

.
Find a b and c of this ellipse 9x²-3y²-36x-6y+12=0
~~~~~~~~~~~~~~~~~~~~~~~~~


It is not an ellipse.

For the given equation, the problem is just solved under this link

https://www.algebra.com/algebra/homework/Quadratic-relations-and-conic-sections/Quadratic-relations-and-conic-sections.faq.question.1125293.html

https://www.algebra.com/algebra/homework/Quadratic-relations-and-conic-sections/Quadratic-relations-and-conic-sections.faq.question.1125293.html



Answer by MathTherapy(10552)


Find a b and c of this ellipse 9x²-3y²-36x-6y+12=0
For a response to this question, see
Quadratic-relations-and-conic-sections/1125292: Find a b and c of this ellipse 9x²-3y²-36x-6y+12=0
OR USE THE FOLLOWING LINK
https://www.algebra.com/algebra/homework/Quadratic-relations-and-conic-sections/Quadratic-relations-and-conic-sections.faq.question.1125292.html